OpenStudy (anonymous):

A cylindrical aquarium has a diameter of 10 ft and a height of 6 ft. What is the exact volume of the aquarium? A. 30π ft³ B. 60π ft³ C. 90π ft³ D. 150π ft³

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@mathmale

geerky42 (geerky42):

Do you know formula of volume of cylinder?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

nevermind I got it IT

geerky42 (geerky42):

\(\large Volume = Base \times height = \pi \times radius^2 \times height\)